OpenStudy (anonymous):

x - 4y=-37 -2x -5y=-53

OpenStudy (anonymous):

Multiply the first equation by 2 so we get 2x-8y=-74 -2x-5y=-53

OpenStudy (anonymous):

Now add them, so we cancel out the x -13y=-74-53

OpenStudy (anonymous):

once you find y, plug into one of the given equation to find x

OpenStudy (anonymous):

idk how to do this

OpenStudy (anonymous):

Okay, what's troubling you?

OpenStudy (anonymous):

all of this

OpenStudy (anonymous):

When you are solving this kind of problem, your goal is to cancel out a variable

OpenStudy (anonymous):

ik but idk howw

OpenStudy (anonymous):

ok so look: -2x+2x=0. Right?

OpenStudy (anonymous):

x - 4y=-37 -2x -5y=-53 Say that you want to cancel out the x, when you add the two equation. Right now it won't cancel. So you have to multiply one equation by some number

OpenStudy (anonymous):

Let's Multiply first equation by 2 we get 2x-8y=-74